Klapperich Lab:Notebook/Lab Meeting Notes/2009/10/20

From OpenWetWare
Jump to navigationJump to search
Owwnotebook icon.png Project name Report.pngMain project page
Resultset previous.pngPrevious entry      Next entryResultset next.png

20 October 2009 Lab Meeting

  • Attending:
  • Missing:
  • Presentation:Dr. Tushar Bansal, post doc applicant. 2pm, please BE ON TIME.

‡ Announcements

  • Oakridge deadline is 1 Feb 2010.
  • PAPERS Drafted before MicroTAS
  • Posters for MicroTAS

‡ Flu R01:Integration

  • New design - address evaporation loss, Reagent delivery and storage solution. Consider marginating flows?

† Sample Concentration (Lead: Jane, Team: Jaephil)

  • MDCK Cells are not propagating. Try troubleshooting efforts. Sonali will do on Monday.
  • Cassidy needs to see plaque assay again - so when cells are up, she needs training.
  • Up to high 10^7 for positive control. Silver substrate no signal.
  • Jane working on the cell lysate control.
  • Main loss is at the outlet/sample collection: tangential filtration design in drawing stage.
  • Contacted EC Shaw about rubber stamp mold for new design for evaporation.
  • Set up COMSOL and StarCD, look for governing equations for simulation of evaporation.
  • Committee meeting setup

† SPE Column Optimization for RNA. (Lead: Sonali, Team: Hussam, Jessie, Brendan)

  • Design three straight channels. A: std channel we already use. B: widest channel possible, C: something in between.Send out masks by 10/21, functional chips to be delivered by 10/26. [Update] functional chips ready by 10/21-22, Experimental results of channels to be delivered on 10/27 [Hussam/Jessie]
  • STD recipe, STD silica beads, just A,B, and C.
  • controls for these: empties, non-silica monolith and full silica monolith, std. recipe.
  • Load DNA in a straw of 5,50,100ul SPE, quick test of hypothesis: more SPE = more DNA, To be repeated with Alex and Mark
  • Look into modifying Alex's pneumatic setup for the PATH guys. Instead of building the straw array over here NOTE: Sean
  • PLanning net silica per channel experiments. with 700nm, 3x possible. Deliver around 11/5/09.

  • New virus prep (JC, 6/11/09), getting Ct's of 27 for 1:16 dilution and 28 for 1:64 dilution. Use more diluted sample for future optimization experiments? Channels were 700nm silica/1X. pfu/ml is 10^6.
  • Tried 2 more chips with 3X 0.15um Silica. Chips are breaking b/c of no crosslinking. SPE was coming out even at 2ml/hr flow rate.
  • Running no silica and no SPE chips with virus. (JC/RNA)

† PCR - CMI (Lead: Qingqing)

  • QQ will work on the initial integration steps of SPE + RT (reservoir)+ PCR.
 * test liquids to fill the empty channel(Mineral oil,Water)
 * Deep reservoir design
 - deformation during bonding
 - fabrication difficulty
 - Fluidic control difficulty
 * third generation design to solve these problem
 - the mold and the chip has been made, and will be tested this week.
  • PCR of C.Difficile DNA
 - ATCC DNA and Primer pairs for both toxin A and B has been ordered.
 - real-time PCR has been done on ABI, waiting for the result.
  • PCR2 Paper formatted for LOAC this week.
  • CMK: PCR 1 draft. Analytical Chem. MCK is running more simulations.

† HDA (Lead: Jaephil, Team:Sonali)

  • Start planning R01 for Submission on 2/5/10. MM, JD, CMK.
  • JD will train QQ how to make HDA chips. This may entail moving from cutter plotter to embossing steps. JD and QQ to meet ASAP to plan the tech transfer.
  • Paper submitted 10/6.

‡C. diff Project (Cathie, Sonali, Satish Singh, Lisa J., His post doc )

  • Sonali to train Lisa on SPE.
  • QQ to run test PCR on chip with genomic DNA and Toxin B primers.

‡ Coulter Flu Fraunhofer Project (Lead: Sonali, Team: Sonali, Jessie, Cathie, CMI Folks, Qingqing)

  • Meeting 9am Monday (every other week)

‡ Agilent Automated Sample Preparation (Lead: Alex)

  • Paul finishes work on nozzles on Wednesday. Then first testrun and rewriting of software for "HotDog"-Run on Machine.
  • First Draft of Paper (Lysis of Yeast Cells) ready. Work ongoing on "Conclusion"-Section of Paper.

‡ COBRA (Lead: Jaephil, Team: Cathie, Jane for virus only, PHO folks)

  • Metal piece modification for the integration - Parts distored and is begin fixed.
  • paper 1: Evap with Sol Gel substrate. Not integrated. MSSA, E coli.

‡ Biointerfaces group (Lead: MinCheol, Team: Cathie, MCK, Wong and Meller folks)

  • New experiments happening now.
  • Cathie will submit paper inquiry.

‡ CIMIT- Sepsis (Lead:Cathie, Team:TBA)

  • IRB approved.
  • Cathie submitted the companion BU IRB form for exemption. Revised, still still waiting.

‡ PATH Grant (Lead:Cathie, Team:Frank, Sean, Mark, Jake Trueb (ME), Suma)

  • Phone meeting with PATH this week. 10/13.
  • UGs purchasing parts for straw machine for 709.
  • UGs purchasing reagents for running extraction experiments and doing PCR.